# Probabilistic Analysis of Copyright Disputes and Generative AI Safety

### v4

@@ 4.1. Evidentiary Considerations in Copyright Disputes Involving AI-Generated Works
When generative AI is used to create content, it is typically offered as a service by the AI developer and used by a user. In such cases, both the user and the AI developer may be considered direct or indirect infringers if the generated output is substantially similar to an original work, depending on the specific circumstances. Copyright infringement may occur even without the user or developer intending to replicate the copyrighted material or to create something similar (Lee et al., 2024).
Due to the uninterpretability of generative models, it can be challenging to provide direct evidence of factual copying for AI-generated content. In most cases, access and similarity will likely serve as the primary evidence of infringement, much like in traditional copyright disputes (Lee et al., 2024). While the creative processes of generative models differ markedly from human creativity, it appears that similar evidentiary principles are likely to apply even when generative AI is involved in the creative processes. At a minimum, it is generally recognized that copying, even for generative AI outputs, requires some form of access to the original work, which may be demonstrated by its inclusion in the AI model’s training data (Zirpoli, 2023; Lee et al., 2024; Elkin-Koren et al., 2024). Moreover, studies have shown that exposure to certain types of content during training leads to a higher likelihood of these models reproducing or closely imitating that content (Carlini et al., 2021, 2023; Hacohen et al., 2024), suggesting that inclusion of specific data correlates with a greater probability of producing similar outputs through copying. This echoes the evidentiary framework applicable to traditional copyright cases, where greater similarity between works strengthens the inference of access and copying.
A distinctive feature of generative models, setting them apart from human artists and traditional creative tools, is their capacity to be trained on vast and diverse datasets. Training datasets are often compiled by scraping content from the Internet, including copyrighted material (Lee et al., 2024). Furthermore, direct inclusion of copyrighted material in the training data is not even necessary for latent diffusion models to produce images similar to copyrighted works, as these models can reproduce content from other images that retain similar latent information (Lu et al., 2024), which exacerbates the problem. These characteristics suggest a high likelihood that generative models have what can be construed as “access” to copyrighted works. As a result, knowing that generative AI was used to create the defendant’s work would likely lead to a stronger belief in the presence of access. For example, consider a scenario where the original work was posted online, and the defendant’s work was created by an AI model trained on data scraped from the Internet. Even in the Internet age, simply posting a work online does not automatically establish access by the defendant. Courts require more than just online availability; factors like the work’s visibility or significant dissemination are needed to strengthen the inference of access. [^8] See generally Cooley v. Target Corp., No. 20-2152 (D. Minn. Sept. 28, 2022). The use of a generative model to produce the defendant’s work can alternatively support the finding of access through the possibility that the model accessed the original work, given the model’s extensive potential to access vast datasets and the common practice of training models on diverse web-scraped data.
4.2. NAF as a Framework for Copyright Safety
The NAF condition (Vyas et al., 2023) ensures that the generation probability of a generative model is closely aligned with that of a version of the model that has not had access to the copyrighted material. In a canonical form (Vyas et al., 2023; Elkin-Koren et al., 2024), the NAF condition can be understood to require the following:
$P(Z=z|Access=1)\leq e^{\epsilon}P(Z=z|Access=0)$
for any $z$, where $Z$ is a random variable representing the output of the generative model, $Access$ indicates whether the model had access to a copyrighted work, and parameter $\epsilon>0$ determines the strength of theoretical guarantee (the smaller the value of $\epsilon$, the closer the output distribution of the model must be to that of a model trained without access). The output distribution of the AI model must satisfy this inequality with respect to any copyrighted work in the training data. The NAF condition with a specific parameter value $\epsilon$ will be referred to as $\epsilon$-NAF herein. One should note that this formulation is idealized, and in reality, the existence or nonexistence of access may not be clear-cut, as copyright-protected elements can be widely distributed and repeated across training data in ways that make them inseparable from other data.
@@ 4.2. NAF as a Framework for Copyright Safety
While the authors suggest that this criterion is related to the concept of substantial similarity in copyright law, they leave open the question of whether this approach could function as a valid defense in court (Vyas et al., 2023). Although the NAF condition has been acknowledged as at least partially effective in reducing the risks of infringement (Elkin-Koren et al., 2024), it has also faced skepticism for not fully aligning with the principles of copyright law (Lee et al., 2024). Here, probabilistic formalism is employed as an aid for precise analysis to reconcile diverging views.
@@
One straightforward way the NAF condition contributes to copyright safety is by addressing the prospective problem of how likely the model is to generate a copyright-infringing output. It can be concluded that this probability is low, given that (i) the model satisfies $\epsilon$-NAF with a small $\epsilon$ and (ii) the model is unlikely to produce outputs highly similar to copyrighted works if it had been trained without access. This trivially follows from standard probabilistic reasoning by viewing the substantial similarity of a generated work to a fixed copyrighted work as a function of the generated work. Specifically, letting $Substantial(Z)$ be a binary function of random variable $Z$ that takes value 1 if the similarity of an output $z$ to the copyrighted work reaches the threshold corresponding to substantial similarity, $\epsilon$-NAF ensures that:
$\begin{split}&P(Substantial(Z)=1|Access=1)\\ &\leq e^{\epsilon}P(Substantial(Z)=1|Access=0).\end{split}$
Due to this inequality, given that $P(Substantial(Z)=1|Access=0)$ is small and $e^{\epsilon}$ is close to 1, corresponding to a small value of $\epsilon$, $P(Substantial(Z)=1|Access=1)$ must also be small. Consequently, the model is unlikely to generate content that is substantially similar to the copyrighted work, and thus unlikely to infringe on its copyright, as its generation would not satisfy the improper appropriation prong.
@@
The focus now shifts to examining whether and how the NAF condition affects the determination of the other prong, factual copying, within the context of the retrospective problem. On one hand, there seems to be little justification for expecting the NAF condition to meaningfully influence the finding of copying, given both the presence of access and a fixed level of similarity in the probative sense. Specifically, the theoretical guarantee of the NAF condition do not directly impose constraints on factual copying; instead, it merely dictates a probabilistic relationship between access and generated outputs. Furthermore, as observed in (Lee et al., 2024), the assertion that a similar output could have been produced by copying from another source does not constitute a valid defense under copyright law.
On the other hand, the NAF condition can impact the inference of access based on similarity, as it influences the observable distribution of similarity between generated content and copyrighted works. Random variables as defined in Definition 4.1 is considered to analyze the implications.
@@
$A_{M}\in\{0,1\}$, $S\in\{1,2,\cdots,n\}$, and $EA_{M}\in\{1,2,\cdots,m\}$ are random variables that satisfy:
@@
In the context of AI generation, access can refer to two distinct concepts: access by the defendant (as in traditional cases) and access by the AI model. To clarify that access by the model is at issue, the variable for access, $A_{M}$, and the corresponding evidence, $EA_{M}$, are denoted with the subscript $M$. It is of interest to see how the NAF condition (Definition 4.1 (1)) affects the inference of access given similarity and other relevant evidence, or in other words, how it affects the value of $P(A_{M}=1|S=i,\,EA_{M}=k)$. To study this, some degree of belief, $\delta_{k}$, about the likelihood of access by the AI model given other evidence without the knowledge of the level of similarity is assumed (Definition 4.1 (2)). Analogous to $EA$ in Definition 3.1, $EA_{M}$ directly impacts only the inference of access (Definition 4.1 (3)).
Given this setting, Proposition 4.2 on the effectiveness of the NAF condition can be obtained.
@@
For any $\epsilon>0$ and $(i,k)$, $P(A_{M}=1|S=i,\,EA_{M}=k)\leq\Gamma(\epsilon,\delta_{k})$ where the random variables are defined as in Definition 4.1 and $\Gamma(\epsilon,\delta):=e^{\epsilon}\delta/(1+\delta(e^{\epsilon}-1))$. $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$, and satisfies $\delta\leq\Gamma(\epsilon,\delta)$ and $\lim_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\delta$.
@@
The proposition shows that the NAF condition indeed imposes an upper bound on the probability of finding the presence of access to the original work by the model. A lower value of $\epsilon$ enhances the effectiveness of this upper bound, and, in the limit of $\epsilon\rightarrow 0$, the evidentiary value of similarity in inferring access vanishes. In other words, similarity fails to serve as meaningful additional evidence of access, with the value of $P(A_{M}=1|S=i,\,EA_{M}=k)$ converging to $\delta_{k}=P(A_{M}=1|EA_{M}=k)$ for all $i$.
However, the way in which the NAF condition operates in this context is ethically debatable. The capacity to generate content with comparable likelihood—regardless of whether access occurred—impedes the determination of whether a work was produced with access based on the generation outcome. Intuitively, since a generative model satisfying the NAF condition is likely to produce a work with the same level of similarity even without access to the copyrighted work, the inferential value of similarity in establishing access by the AI model is reduced. It is questionable whether such an approach is justified, especially considering the demand for transparency in training data as advocated by researchers (Sag, 2023; Zhang et al., 2024) and reflected in the EU AI Act, [^9] Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence and amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act), Official Journal of the European Union, L 1689, 12 July 2024, pp. 1–144. which imposes transparency requirements on developers of certain AI systems, mandating that they make available summaries of the training data used. Furthermore, since $\Gamma$ is bounded below by the degree of belief $\delta_{k}$ that the model had access to the specific original work in question, given available evidence of access other than similarity, the upper bound on the degree of belief about access becomes increasingly insignificant as stronger evidence of access to the specific original work by the model is presented (e.g., unequivocal evidence of the work’s inclusion in its training data) or as it becomes common practice to include broader sets of copyrighted materials in the training of AI models in general, both of which can lead to higher values of $\delta_{k}$.
5. Discussion
Probabilistic analysis has been shown to enhance the understanding of copyright jurisprudence and its interaction with emerging technologies. In particular, a mathematical proof demonstrated that the controversial inverse ratio rule holds, provided it is appropriately defined. Furthermore, the NAF condition was examined as a mitigation strategy for the copyright infringement risks posed by generative AI, revealing ethical concerns and limitations in its efficacy when addressing the retrospective problem adjudicators face in court, where they must determine whether infringement has occurred based on available evidence.
@@ 5. Discussion
There are certain limitations to the analysis presented in this paper. While the analysis is based on a probabilistic view of legal fact-finding, this perspective has faced criticisms (Tribe, 1971; Allen and Stein, 2013), although it continues to receive support (Cheng, 2013; Burtis et al., 2017). The validity of probabilistic approaches in legal fact-finding in general is widely debated in legal literature and falls beyond the scope of this paper, except for the relation between the results concerning the inverse ratio rule and the conjunction problem. Furthermore, the analysis abstracts from certain details involved in AI generation, such as the content of user-provided prompts, which may influence the identification of direct or indirect infringers (Lee et al., 2024) and be important in actual cases.
Several related directions for future work remain to be explored. Although this paper has paid particular attention to the inverse ratio rule—especially compatible with probabilistic reasoning due to its direct connection to the assessment of evidence—the presented approach has broader potential applications within copyright jurisprudence, as many copyright principles can be understood through the lens of evidentiary rules (Lichtman, 2003). Furthermore, while the analysis has addressed the NAF condition, future research should focus on developing alternative copyright risk mitigation strategies beyond the NAF condition and conducting a more comprehensive analysis of how generative AI technology interacts with copyright law.
@@ Appendix A The Inverse Ratio Rule with Strict Inequality
Proposition 3.2 was proved by assuming only very general relationships among the $\alpha_{i}$’s as well as among the $\beta_{i,j}$’s. It does not specify, for example, how the value of $\beta_{i,j}$ changes when one index is increased while the other is decreased simultaneously, nor does it address the relative magnitude of the changes in the values of $\alpha_{i}$ and $\beta_{i,j}$. Proposition A.1 demonstrates that, under certain assumptions regarding these details, the inverse ratio rule holds with strict inequality.
The inequality $\alpha_{\mu^{\prime}}/\alpha_{\mu^{\prime}-1}\leq\beta_{\mu^{\prime}-1,j}/\beta_{\mu^{\prime},j^{\prime}}$ assumed in the first part of Proposition A.1 roughly states that the increase in the degree of belief in access resulting from stronger evidence of access (excluding similarity) is large relative to the changes in the degrees of belief in both access and copying that arise from different levels of similarity. On the other hand, the inequality $\beta_{i^{\prime},\kappa^{\prime}}/\beta_{i,\kappa^{\prime}-1}\leq\alpha_{i}/\alpha_{i^{\prime}}$ in the second part of the proposition roughly means that the increase in the degree of belief in copying due to a higher level of similarity is substantial relative to the possible increase (or decrease, in which case the inequality is trivially satisfied) in the degree of belief in access with stronger evidence of access paired with a showing of a lower level of similarity.
Importantly, these relations are local in the sense that they only have to hold for some specific indices for the proposition to apply, and it is not necessary for them to be universally true across all indices $i$ and $j$. It is quite plausible, therefore, that the proposition apply in some actual instances.
@@
Assuming that the random variables are defined as in Definition 3.1, the following strong inverse ratio rule holds for any $\lambda\in(0,1)$:
  • Suppose $j>j^{\prime}$ and $\mu^{\prime}>1$ where $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$. Then, $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}<\mu^{\prime}$ if $\alpha_{\mu^{\prime}}/\alpha_{\mu^{\prime}-1}\leq\beta_{\mu^{\prime}-1,j}/\beta_{\mu^{\prime},j^{\prime}}$.
@@
  • Suppose $i>i^{\prime}$ and $\kappa^{\prime}>1$ where $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$. Then, $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}<\kappa^{\prime}$ if $\beta_{i^{\prime},\kappa^{\prime}}/\beta_{i,\kappa^{\prime}-1}\leq\alpha_{i}/\alpha_{i^{\prime}}$.
Appendix B Relation to the Conjunction Problem
The conjunction problem is frequently cited to criticize probabilistic approaches to legal fact-finding (Pardo and Allen, 2008). It refers to the issue that when a claim consists of multiple elements that must each be established by evidence, the probability of the conjunction of these elements may not reach a required threshold $\lambda$, even if the probability of each individual element exceeds $\lambda$. More concretely, suppose there are two events, $E_{1}$ and $E_{2}$, that need to be established to make a certain claim. The probability that both occurs is the product of the probabilities of the two elements, $P(E_{1}\cap E_{2})=P(E_{1}|E_{2})P(E_{2})$, which is smaller than the probability of either individual element. Thus, even if both elements have been shown to have probabilities higher than $\lambda$ individually, $P(E_{1}|E_{2}),P(E_{2})>\lambda$, it can be the case that the probability of both requirements being satisfied simultaneously is below the threshold, $P(E_{1}\cap E_{2})<\lambda$. A solution offered by the probabilistic camp is to require that the probability of the conjunction, in addition to the individual probability of each element, exceeds the threshold in order to make an affirmative finding (Nance, 1986; Friedman, 1992). However, critics object to this solution for its counterintuitiveness (Pardo and Allen, 2008).
Such objection may appear to affect the results in this paper regarding the inverse ratio rule, which relies on how the probability of copying being a product of two terms: the probability of copying given the presence of access and the probability of access. However, this is not the case because the situation here differs from those considered in the context of the conjunction problem. Under the Arnstein test, it is not required to prove access to the original work as a separate element from copying. Instead, the plaintiff merely needs to show that the probability of copying reaches the threshold, $P(C=1|S=i,\,EA=j)>\lambda$. As shown in the proof of Proposition 3.1, due to the logical relationship between copying and access (Definition 3.1 (1)), this probability can be expressed as the product of two terms, the probability of copying and that of access, leading to:
$\begin{split}&P(C=1|S=i,\,EA=j)>\lambda\\ &\Leftrightarrow P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)>\lambda.\end{split}$
This product arises merely as a consequence of the logical relationship between copying and access, rather than from adopting the controversial solution that necessitates the probability of the conjunction reaching the threshold, which can be formally described as:
$P(C=1,\,A=1|S=i,\,EA=j)>\lambda.$
Therefore, the argument presented here is not directly influenced by discussions surrounding the conjunction problem.
Appendix C Proofs of Propositions
C.1. Proof of Proposition 3.2

### v5

@@ 3. Near Access-Freeness as a Framework for Copyright Safety of Generative AI
When generative AI is used to create content, it is typically offered as a service by the AI developer and used by a user. In such cases, both the user and the AI developer may be considered direct or indirect infringers if the generated output is substantially similar to an original work, depending on the specific circumstances. Copyright infringement may occur even without the user or developer intending to replicate the copyrighted material or to create something similar (Lee et al., 2024). Due to the uninterpretability of generative models, it can be challenging to provide direct evidence of factual copying for AI-generated content. In most cases, access and similarity will likely serve as the primary evidence of infringement, much like in traditional copyright disputes (Lee et al., 2024). The inclusion of a work in a model’s training corpus is often viewed as sufficient evidence of the access element required to establish copying (Lee et al., 2024; Elkin-Koren et al., 2024).
Near Access-Freeness (Vyas et al., 2023), which aims to mitigate copyright infringement risks, ensures that the generation probability of a generative model is closely aligned with that of a version of the model that has not had access to copyrighted data. In a canonical form (Vyas et al., 2023; Elkin-Koren et al., 2024), the NAF condition can be understood to impose the inequality $P(Z=z|Access=1)\leq e^{\epsilon}P(Z=z|Access=0)$ for any $z$, where $Z$ is a random variable representing the output of the generative model, $Access$ indicates whether the model had access to a copyrighted work, and parameter $\epsilon>0$ determines the strength of theoretical guarantee. The output distribution of the AI model must satisfy this inequality with respect to any copyrighted work in the training data. In words, if a model meets the NAF condition with some $\epsilon$, having access to the original work during training can raise the probability of generating any particular output by at most a factor of $e^{\epsilon}$. For instance, a model trained on Snoopy images is at most $e^{\epsilon}$ times more likely to produce any image—including one that resembles Snoopy—than a version of the model that never had access to Snoopy images in training. The NAF condition with a specific parameter value $\epsilon$ will be referred to as $\epsilon$-NAF herein.
@@
The authors who proposed the NAF condition leave open the question of whether this approach could function as a valid defense in court (Vyas et al., 2023). Although the NAF condition has been acknowledged to be at least partially effective in reducing the risks of infringement (Elkin-Koren et al., 2024), it has also faced skepticism for not fully aligning with the principles of copyright law (Lee et al., 2024). Here, a probabilistic formalism is employed as an aid for precise analysis to reconcile divergent views.
@@
One straightforward way the NAF condition contributes to copyright safety is by addressing the prospective problem of how likely the model is to generate a copyright-infringing output. It can be concluded that this probability is low, given that (i) the model satisfies $\epsilon$-NAF with a small $\epsilon$ and (ii) the model is unlikely to produce outputs highly similar to copyrighted works if it had been trained without access. This trivially follows from standard probabilistic reasoning by viewing the substantial similarity of a generated work to a fixed copyrighted work as a function of the generated work. Specifically, letting $Substantial(Z)$ be a binary function of random variable $Z$ that takes value 1 if the similarity of an output $z$ to the copyrighted work reaches the threshold corresponding to substantial similarity, $\epsilon$-NAF ensures that: $P(Substantial(Z)=1|Access=1)\leq e^{\epsilon}P(Substantial(Z)=1|Access=0)$. Due to this inequality, given that $P(Substantial(Z)=1|Access=0)$ is small and $e^{\epsilon}$ is close to 1, corresponding to a small value of $\epsilon$, $P(Substantial(Z)=1|Access=1)$ must also be small. Consequently, the model is unlikely to produce substantially similar content or infringe copyright, as it would not meet the improper appropriation prong.
@@
The focus now shifts to examining whether and how the NAF condition affects the determination of the other prong, factual copying, within the context of the retrospective problem. On one hand, there seems to be little justification for expecting the NAF condition to meaningfully influence the finding of copying, given both the presence of access and a fixed level of similarity in the probative sense. Specifically, the theoretical guarantee of the NAF condition does not directly impose constraints on factual copying; instead, it merely dictates a probabilistic relationship between access and generated outputs. Furthermore, as observed in (Lee et al., 2024), the assertion that a similar output could have been produced by copying from another source does not constitute a valid defense under copyright law.
On the other hand, the NAF condition can impact the inference of access based on similarity, as it influences the observable distribution of similarity between generated content and copyrighted works. This becomes evident through a formal analysis employing the random variables as defined in Definition 3.1.
@@
$A_{M}\in\{0,1\}$ and $EA_{M}\in\{1,2,\cdots,m\}$ are random variables that satisfy:
@@
In the context of AI generation, access can refer to access by the user or access by the AI model. To clarify that access by the model is at issue, the variable for access, $A_{M}$, and the corresponding evidence, $EA_{M}$, are denoted with the subscript $M$. The NAF condition (Definition 3.1 (1)) and some degree of belief, $\delta_{k}$, about the likelihood of access by the AI model given available evidence (other than similarity) of access (Definition 3.1 (2)) are assumed. Analogous to $EA$ in Definition 2.1, $EA_{M}$ directly impacts only the inference of access (Definition 3.1 (3)).
@@
For any $\epsilon>0$ and $(i,k)$, $P(A_{M}=1|S=i,\,EA_{M}=k)\leq\Gamma(\epsilon,\delta_{k})$ where $\Gamma(\epsilon,\delta):=\frac{e^{\epsilon}\delta}{1+\delta(e^{\epsilon}-1)}$. $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$, and satisfies $\delta\leq\Gamma(\epsilon,\delta)$ and $\lim\limits_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\delta$.
@@
Proposition 3.2 shows that the NAF condition indeed imposes an upper bound on the probability of finding the presence of access to the original work by the model. A lower value of $\epsilon$ enhances the effectiveness of this upper bound, and, in the limit of $\epsilon\rightarrow 0$, the evidentiary value of similarity in inferring access vanishes. In other words, similarity fails to serve as meaningful additional evidence of access, with the value of $P(A_{M}=1|S=i,\,EA_{M}=k)$ converging to $\delta_{k}=P(A_{M}=1|EA_{M}=k)$ for all $i$.
However, the way in which the NAF condition operates in this context is ethically debatable. Intuitively, when the NAF condition is satisfied with small $\epsilon$, training with access to the original work does not increase the model’s probability of generating any particular output much (not more than a multiplicative factor $e^{\epsilon}$) compared to training without access. This, in turn, ensures that the model is not significantly more likely to produce outputs similar to the original work relative to a version of the model that did not have access. Hence, one cannot strongly infer that the model must have been trained on the original work just because it produces a similar output. However, this approach may contradict the demand for transparency in training data, as advocated by researchers (Sag, 2023; Yoo, 2024). Furthermore, $\Gamma$ is bounded below by the degree of belief $\delta_{k}$ that the model had access to the specific original work in question, given available evidence of access other than similarity. The upper bound on the degree of belief about access becomes less significant as the baseline belief $\delta_{k}$ increases, either through stronger evidence of access to the original work by the model (e.g., unequivocal evidence of its inclusion in the training data) or through the increasingly common practice of training AI models on broader sets of copyrighted materials.
4. Concluding Remarks
@@
Probabilistic analysis has been shown to enhance the understanding of copyright jurisprudence and its interaction with generative AI. However, there are certain limitations to the analysis presented in this paper. While the analysis is based on a probabilistic view of legal fact-finding, this perspective has faced sustained criticism (Tribe, 1971; Allen and Stein, 2013); evaluating the validity of probabilistic approaches in legal fact-finding in general falls beyond the scope of this paper. Furthermore, the analysis abstracts from certain details involved in AI generation, such as the content of user-provided prompts, which may influence the identification of direct or indirect infringers (Lee et al., 2024).
Several future directions remain to be explored. Firstly, the presented approach has wider potential applications within copyright jurisprudence, where many legal doctrines can be understood from evidentiary perspectives (Lichtman, 2003). In addition, a variety of formal and computational methodologies (Verheij et al., 2015; Halpern, 2017) for evidentiary and probabilistic reasoning beyond the simple framework used here could be applied to further refine and operationalize the probabilistic approach. Lastly, future research should analyze and develop copyright risk mitigation strategies alternative to the NAF condition.
@@ Appendix A The Inverse Ratio Rule with Strict Inequality
The inverse ratio rule holds with strict inequalities under moderate additional assumptions (Proposition A.1).
@@
The following strong inverse ratio rule holds for any $\lambda\in(0,1)$: Suppose $j>j^{\prime}$ and $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}>1$. If $\tfrac{\alpha_{\mu^{\prime}}}{\alpha_{\mu^{\prime}-1}}\leq\tfrac{\beta_{\mu^{\prime}-1,j}}{\beta_{\mu^{\prime},j^{\prime}}}$, then $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}<\mu^{\prime}$. Similarly, suppose $i>i^{\prime}$ and $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}>1$. If $\tfrac{\beta_{i^{\prime},\kappa^{\prime}}}{\beta_{i,\kappa^{\prime}-1}}\leq\tfrac{\alpha_{i}}{\alpha_{i^{\prime}}}$, then $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}<\kappa^{\prime}$.
Appendix B Proofs
Proof of Proposition 2.2.